# Projected Gradient Descent Method for Tropical Principal Component Analysis over Tree Space

### v1

@@ Abstract.
In 2019, Yoshida et al. developed tropical Principal Component Analysis (PCA), that is, an analogue of the classical PCA in the setting of tropical geometry and applied it to visualize a set of gene trees over a space of phylogenetic trees which is an union of lower dimensional polyhedral cones in an Euclidean space with its dimension $m(m-1)/2$ where $m$ is the number of leaves. In this paper, we introduce a projected gradient descent method to estimate the tropical principal polytope over the space of phylogenetic trees and we apply it to apicomplexa dataset. With computational experiment against MCMC samplers, we show that our projected gradient descent works very well.
@@ 6. Conclusion
In this short paper, we introduce a novel method to approximate the best-fit tropical polytope to explain a sample of gene trees. We show that this gradient method reduces the objective function with appropriate learning rate and it works very well from the computational experiment. In an experiment, we implement a learning rate that decreases with each iteration, but it is not clear how we schedule learning rates for this problem.
We implement our method in R and the source code is available by requesting the author.

### v2

@@ Abstract.
In 2019, Yoshida et al. developed tropical Principal Component Analysis (PCA), that is, an analogue of the classical PCA in the setting of tropical geometry and applied it to visualize a set of gene trees over a space of phylogenetic trees which is an union of lower dimensional polyhedral cones in an Euclidean space with its dimension $m(m-1)/2$ where $m$ is the number of leaves. In this paper, we introduce a projected gradient descent method to estimate the tropical principal polytope over the space of phylogenetic trees and we apply it to apicomplexa dataset. With computational experiment against MCMC samplers, we show that our projected gradient descent has a lower sum of tropical distances between observations and their projections on an estimated best-fit tropical polytope compared with the Markov Chain Monte Carlo (MCMC) approach proposed by Page et al.
@@ 5. Computational Experiments
5.1. Simulated Dataset
In this experiments, we generate gene trees from the multispecies coalescent model (MCM) with a given species tree via the software Mesquite [11]. We fixed the effective population size $N_{e}=100,000$ and varied $R=\frac{SD}{N_{e}}$ where $SD$ is the species depth which is the number of generations from the common ancestor (the root of the tree) to its leaves.
In this experiment, we simulate $1000$ trees for each $R=0.25,0.5,1.0,2.0,5.0,10.0$ and we set iteration to be 100 with a MCMC method as well as our projected gradient method. We recompute a MCMC technique and our projected gradient method 10 times for each $R$ so that we can capture variation of each sum of tropical distance between observations and their projections on an estimated tropical polytope (sum of ”magnitude of errors” (SE) in terms of $d_{\rm tr}$), i.e., an estimated optimal value of the linear programming problem in Problem 1. The results are shown in Figure 2. For the scheduling of the learning rate we start $\lambda=0.001$. Then each iteration we multiply the learning rate by $0.999$. Note that this learning rate is not optimized. Therefore, we could investigate what is the optimal learning rates. We set the stopping criteria if the iteration reaches to the maximum iteration which we preset. Each iteration should take $O(m^{2}n)$, where $n$ is the sample size.
Figure 2. Side-by-Side Boxplots for SE for each method and each different ratio. We repeat computation 10 times for each ration and each method.
5.2. Empirical Dataset
@@
We set the maximum iteration to be 100 for our projected gradient method and for a MCMC method, we set the maximum iterations to be 1000. For the scheduling of the learning rate we start $\lambda=0.01$. Then each iteration we multiply the learning rate by $0.999$. Note that this learning rate and scheduling are not optimized.
@@ 6. Conclusion
In this short paper, we introduce a novel method to approximate the best-fit tropical polytope to explain a sample of gene trees. We show that this gradient method reduces the objective function with appropriate learning rate and it has a lower sum of tropical distances between observations and their projections on an estimated best-fit tropical polytope compared with the MCMC approach proposed by Page et al. [12] from the computational experiment. In an experiment, we implement a learning rate that decreases with each iteration, but it is not clear how we schedule learning rates for this problem.
We implement our method in R and the source code is available at http://polytopes.net/Tropical_Gradient2.zip.